# Multi-FAct: Assessing Multilingual LLMs’ Multi-Regional Knowledge using FActScore

### 28 Feb 2024 v1

@@ A.2.2 Evaluation of Multi-FAct Pipeline
The basis of our pipeline relies crucially on the GPT3.5-based translation (Step 3 in figure 2). It is imperative to see whether this translation step hurts the estimation of FActScore or not.
@@
The basis of our evaluation method is that the factuality of a text is translation invariant, ie, factual knowledge in a text should not change when it is translated to other languages. To evaluate this, we first take the human-annotated examples provided in original FActScore , and translate them into four languages666one language from one continent using GPT4. We use GPT4 as previous studiesJiao et al. (2023) revealed GPT4 to be a good translator and better than GPT3.5 and commercially available machine translators. Native speakers of a subset of our languages also confirmed this.
@@
We then applied our multilingual FActScore pipeline (Figure 2) on this GPT4 translated text. The premise is that if our pipeline (which involves GPT3.5 translation [^7] We did not use GPT4 translation in Multi-FAct pipeline to reduce costs) provides similar FActScore as the original English annotated data, it might provide good evidence that having a translation step in our pipeline doesn’t hurt FActScore estimation significantly.
Figure 9: Correlation of FActScore after on translated human evaluation data.
The results are provided in Table 4. We see that running Multi-FAct on non-English data does not significantly degrade FActScore evaluation.

### 01 Mar 2024 v2

@@ A.2.2 Evaluation of Multi-FAct Pipeline
The basis of our pipeline relies crucially on the GPT3.5-based translation (Step 3 in figure 2). It is imperative to see whether this translation step hurts the automated estimation of FActScore in a multilingual setting.
@@
The basis of our evaluation method is that the factuality of a text is translation invariant, ie, factual knowledge in a text should not change when it is translated to other languages. To evaluate this, we first take the human-annotated examples provided in original FActScore , and translate them into four languages using GPT4. We use GPT4 as previous studies Jiao et al. (2023) revealed GPT4 to be a good translator and better than GPT3.5 and commercially available machine translators. Manual inspection by native speakers of a subset of our languages also confirmed this.
@@
Figure 9: Correlation of FActScore evaluated on translated human-annotated data.
We then applied our multilingual FActScore pipeline (Figure 2) on this GPT4 translated text. The premise is that if our pipeline (which involves GPT3.5 translation [^6] We did not use GPT4 translation in Multi-FAct pipeline to reduce API costs) provides similar FActScore as the original English annotated data, it might provide good evidence that having a translation step in our pipeline doesn’t hurt FActScore estimation significantly.
The results are provided in Table 4. We see that running Multi-FAct on non-English data does not significantly degrade FActScore estimation.